# Lagrangian extensions and left symmetric structures on the four-dimensional real Lie superalgebras

### v1

@@ Abstract.
Over real numbers, Backhouse classified all four-dimensional Lie superalgebras. From this list, we will investigate those Lie superalgebras that can be obtained as Lagrangian extensions. Moreover, we investigate left-symmetric structures on these Lie superalgebras. Furthermore, except for two of them, they are all Novikov superalgebras.
@@
Quasi-Frobenius Lie (super)algebra, Lagrangian extensions, Left symmetric structures, Novikov structures, Balinsky-Novikov structures.
@@ 1.1. Lagrangian extensions of Lie (super)algebras
The notion of $T^{*}$-extension of Lie algebras was initially introduced by Bordemann [Bor]. A short explanation can be understood as the following. Given a Lie algebra $\mathfrak{h}$ equipped with an $\mathfrak{h}^{*}$-valued bilinear map $\alpha$, a $T^{*}$-extension of $\mathfrak{h}$ is the Lie algebra structure on the vector space $\mathfrak{g}:=\mathfrak{h}\oplus\mathfrak{h}^{*}$ constructed by means of the form $\alpha$ that satisfies a certain cohomological requirement. The Lie algebra $\mathfrak{g}$ naturally admits an invariant non-degenerate symmetric bilinear form. Conversely, Bordemann showed in [Bor] that every finite-dimensional nilpotent Lie algebra $\mathfrak{h}$ with an invariant symmetric non-degenerate bilinear form “is” a suitable $T^{*}$-extension. In particular, if $\mathfrak{h}$ is a real finite-dimensional Lie algebra belonging to a Lie group $G$, then the $T^{*}$-extension of $\mathfrak{h}$ is the Lie algebra of the cotangent bundle $T^{*}G$ of $G$, considered as a Lie group. The terminology and notation are justified by this differential geometric fact.
A superization of this construction was given in [BBB]. In particular, it was shown that every solvable Lie superalgebra with an even symmetric non-degenerate bilinear form is either isomorphic to a $T^{*}$-extension of a certain Lie superalgebra or to an ideal of codimension one of a certain $T^{*}$-extension.
@@
Lie algebras possessing both an invariant non-degenerate symmetric bilinear form and a non-degenerate 2-cocycle of its scalar cohomology, were studied in [BBM] and were called quadratic symplectic Lie algebras. If the ground field is algebraically closed, the authors showed that every quadratic symplectic Lie algebra is a $T^{*}$-extension of a Lie algebra which admits an invertible derivation. Necessary and sufficient conditions on $\mathfrak{h}$ and on the cocycle used in the construction of the $T^{*}$-extension are investigated to ensure that the extended algebra admits a skew-symmetric derivation and, hence, a symplectic structure. As a result, complex symplectic quadratic Lie algebras with dimensions less than or equal to 8 are classified. Some of these Lie algebras correspond to some Lie groups that admit a bi-invariant pseudo-Riemannian metric as well as a left-invariant symplectic form. These Lie groups are nilpotent, and their geometry is very rich, see [BBM, MR] for more details.
The notion of Lagrangian extension was superized in [BM]. The construction is valid for arbitrary Lie superalgebras defined over a field of characteristic not two provided the Lie superalgebra admit a torsion-free flat connection. Moreover, they have observed that there are two ways to perform this extension: they consider either $\mathfrak{g}:=\mathfrak{h}\oplus\mathfrak{h}^{*}$ or $\mathfrak{g}:=\mathfrak{h}\oplus\Pi(\mathfrak{h}^{*})$, where $\Pi$ is the change of parity functor. They call them $T^{*}$-extension and $\Pi T^{*}$-extensions of $\mathfrak{h}$, respectively. In the first case, the non-degenerate form is even, and in the second case, it is odd.
@@
One of the main objectives of this paper is to complete the list of examples among the 4-dimensional Lie superalgebras and explain how they can be obtained as Lagrangian extensions, see Section 4. According to the Backhouse classification, this investigation will classify all four-dimensional real Lie superalgebras that admit (or don’t admit) Lagrangian extensions.
@@ 1.2. Left-symmetric structures
Left-symmetric algebras (also known as Vinberg-Koszul algebras) play an important role in various fields of mathematics and mathematical physics; for instance, rooted tree algebras, operad theory, vertex algebras, convex homogeneous cones, affine manifolds, left-invariant affine structures on Lie groups, to name just a few. For a thorough review, see [B3].
Over the real and complex numbers left-symmetric algebras are closely related to the theory of left-invariant affine structures over Lie groups, see for example [M, GS] (and references therein).
The investigation and classification of left-symmetric structure on Lie (super)algebras is an important subject by its own, as initiated in [S]. The problem, however, seems to be difficult to handle.
For reductive Lie algebras, Baues [Ba] proved that $\mathfrak{gl}(n)$ is the only reductive Lie algebra with one-dimensional center and a simple semisimple ideal which admits left-symmetric algebras over an algebraically closed ground field (see also [B1]). Morover, all left-symmetric algebras on $\mathfrak{gl}(n)$ were classified.
Simple Lie algebras over complex numbers do not admit left-symmetric structures, see [H]. The existence of such structures is not always guaranteed, even for nilpotent Lie algebras. There are filiform nilpotent Lie groups of dimension $10\leq n\leq 13$ which do not admit any left-invariant affine structure, see [B, B3, BG]. Any filiform nilpotent Lie group of dimension $n\leq 9$ admits a left-invariant affine structure. Nevertheless, left-symmetric algebra structures on both finite-dimensional and infinite-dimensional Lie algebras have been studied extensively, see [Bai, B2, YZ] as well as [CL, KCB, TB], respectively. Left symmetric structures on the Virasoro superalgebra were studied in [KB].
In the super setting, it was proved in [DZ] that $\mathfrak{sl}(m+1|m)$ admits a left-symmetric superalgebra, and a full classification is offered in the case of $\mathfrak{sl}(2|1)$. Moreover, it was shown that $\mathfrak{sl}(m|1)$ does not admit left-symmetric superalgebras for $m\geq 3$.
@@
A second aim of this paper is to provide explicitly a left-symmetric structure on the 4-dimensional real Lie superalgebras. Surprisingly enough, all these Lie superalgebras admit such a structure, given explicitly in Section 5. Furthermore, we show that these structures are Novikov, except for the two Lie superalgebras $(D_{0}^{10})^{1}$ and $(D_{0}^{10})^{2}$. In addition, all these Lie superalgebras admit a Balinsky Novikov structure. Classifying all left-symmetric structures on each of these Lie superalgebras remains an open problem.
@@ 2. Main concepts and definitions
Throughout the text $\mathbb{K}$ is a field of characteristic $p\not=2$.
A Lie superalgebra is a $\mathbf{Z}_{2}$-graded space $\mathfrak{g}=\mathfrak{g}_{\bar{0}}\oplus\mathfrak{g}_{\bar{1}}$ over a field of characteristic $p\not=2$ together with a bilinear binary operator $[\cdot,\cdot]:\mathfrak{g}\times\mathfrak{g}\rightarrow\mathfrak{g}$, called the Lie superbracket that has the following properties: (for any homogeneous $x,y,z\in\mathfrak{g}$)
  1. (i)
    Super anti-commutativity: $[x,y]=-(-1)^{|x||y|}[y,x]$;
  2. (ii)
    Super Jacobi identity: $[x,[y,z]]=[[x,y],z]+(-1)^{|x||y|}[y,[x,z]]$ where $|x|={\bar{0}}$ if $x\in\mathfrak{g}_{\bar{0}}$ and $|x|={\bar{1}}$ if $x\in\mathfrak{g}_{\bar{1}}$;
  3. (iii)
    If $p=3$, one must add $[x,[x,x]]=0$ if $x\in\mathfrak{g}_{\bar{1}}$.
Let $V$ and $W$ be two superspaces defined over an arbitrary field ${\mathbb{K}}$. Let $\omega\in\text{Bil}(V,W)$ be a homogeneous bilinear form. The Gram matrix $B=(B_{ij})$ associated to $\omega$ is given by the formula,
(1) $B_{ij}=(-1)^{|\omega||v_{i}|}\omega(v_{i},v_{j})\text{\leavevmode\nobreak\ \leavevmode\nobreak\ for the basis vectors $v_{i}\in V$.}$
(2) $u(\omega)(u,v)=(-1)^{|v||u|}\omega(v,u)\text{\leavevmode\nobreak\ \leavevmode\nobreak\ for any homogeneous $v\in V$ and $u\in W$.}$
In terms of the Gram matrix $B$ of $\omega$: the form $\omega$ is symmetric if and only if
(3) $u(B)=B,\;\text{ where $u(B)=\left(\begin{array}[]{cc}R^{t}&(-1)^{|\omega|}T^{t}\\[5.69054pt] (-1)^{|\omega|}S^{t}&-U^{t}\end{array}\right),$ for $B=\left(\begin{array}[]{cc}R&S\\[5.69054pt] T&U\end{array}\right)$.}$
Similarly, anti-symmetry of $\omega$ means that $u(B)=-B$.
An even (resp. odd) non-degenerate bilinear form is called orthosymplectic (resp. periplectic).
Proposition 2.1 (Superdimension constraints).
If $V=V_{\bar{0}}\oplus V_{\bar{1}}$ is a finite-dimensional superspace such that $V_{\bar{1}}\not={0}$, equipped with an anti-symmetric non-degenerate bilinear form $\omega$, then:
  1. (i)
    If $p(\omega)=\bar{0}$, then $\mathrm{dim}(V_{\bar{0}})$ is even.
  2. (ii)
    If $p(\omega)=\bar{1}$, then $\mathrm{dim}(V_{\bar{0}})=\mathrm{dim}(V_{\bar{1}})$.
Proof.
See, [BM]
We will denote by $\Pi$ the functor that assigns to $V$ the superspace $\Pi(V)$ defined as the other copy of $V$ with the opposite parity of its homogeneous components. We have:
@@
Let $\omega\in\mathrm{Bil}(\mathfrak{g},\mathfrak{g})$ be an anti-symmetric bilinear form on a Lie superalgebra $\mathfrak{g}$. The form $\omega$ is called closed if it satisfies:
@@
for all $x,y,z\in\mathfrak{g}_{\bar{i}}$. This means that $\omega$ is 2-cocyle with scalar values.
A Lie superalgebra $\mathfrak{g}$ is called quasi-Frobenius if it is equipped with a closed anti-symmetric non-degenerate bilinear form $\omega$. If $\omega$ is even (resp. odd) we call it orthosymplectic (resp. periplectic). We denote such a Lie superalgebra by $(\mathfrak{g},\omega)$. Recently, these Lie superalgebras were studied in [BE, BEM, BM] in the context of Lagrangian and double extensions.
@@ 2.1. Connections on Lie (super)algebras
Let $\nabla:\mathfrak{g}\times\mathfrak{g}\rightarrow\mathfrak{g}$ be a connection on $\mathfrak{g}$. For all $x,y,z\in\mathfrak{g}_{\bar{i}}$, the torsion $T^{\nabla}$ of $\nabla$ is defined as
@@
2.2. Left-Symmetric Lie (super)algebras
@@
These superalgebras are Lie-admissible. The supercommutator $[x,y]:=x\cdot y-(-1)^{|x||y|}y\cdot x$ defines then a Lie superalgebra structure on $L$ and we denote it by ${\mathfrak{g}}(L)$. The resulting Lie superalgebra is called the associated Lie superalgebra of the LSSA. Note that if $L$ admits an LSSA then the even part $L_{\bar{0}}$ also admits one. Associative superalgebras are all LSSAs.
In [H], Helmstetter showed that it is necessary to have $[L,L]\subsetneq L$ if $L$ is a left symmetric algebra defined over the field of complex numbers. Over a field of characteristic $p$, this is not true anymore. Burde showed in [B2] that classical Lie algebras over a field $p>3$ admit an LSS only in case $p$ divides $\mathrm{dim}(L)$. In the super setting and over a field of prime characteristic, there has not been much research done so far.
@@
Additionally, every quasi-Frobenius Lie superalgebra $({\mathfrak{g}},\omega)$ admits a left symmetric structure given by
@@
The converse is not always true. There will be several examples that illustrate that in Section 5.
Some subclasses of LSSAs are important as well. An LSSA with supercommutative right multiplication is called a Novikov superalgebra (cf. [BN, Xu]). Explicitly, a Novikov superalgebra $(L,\cdot)$ is an LSSA with an additional condition:
@@
    Left symmetry and commutativity:
@@
    $\begin{array}[]{ll}(x\cdot y)\cdot z=(x\cdot z)\cdot y&\text{ for all }x,y\in L_{\bar{1}}\text{ and }z\in L_{\bar{0}};\\[5.69054pt] x\cdot(y\cdot z)=(x\cdot y)\cdot z+y\cdot(x\cdot z)&\text{ for all }x\in L_{\bar{0}}\text{ and }y,z\in L_{\bar{1}};\\[5.69054pt] x\cdot(y\cdot z)=(x\cdot y)\cdot z+(x\cdot z)\cdot y&\text{ for all }x,y,z\in L_{\bar{1}}.\\[5.69054pt] \end{array}$
@@
$\begin{array}[]{ll}[x,y]:=x\cdot y-y\cdot x,&\text{ for all }x,y\in L_{\bar{0}};\\[5.69054pt] [x,y]:=x\cdot y-\frac{1}{2}y\cdot x=-[y,x],&\text{ for all }x\in L_{\bar{0}},\text{ and }y\in L_{\bar{1}}\\[5.69054pt] [x,y]:=x\cdot y&\text{ for all }x,y\in L_{\bar{1}}.\end{array}$

### v2

# Lagrangian extensions and left-symmetric structures on the four-dimensional real Lie superalgebras

@@ Abstract.
Over real numbers, Backhouse classified all four-dimensional Lie superalgebras. From this list, we will investigate those Lie superalgebras that can be obtained as Lagrangian extensions. Moreover, we investigate left-symmetric structures on these Lie superalgebras. Furthermore, except for two of them, we show that they are all Novikov superalgebras.
@@
Quasi-Frobenius Lie (super)algebra, Lagrangian extensions, Left-symmetric structures, Novikov structures, Balinsky-Novikov structures.
@@ 1.1. Lagrangian extensions of Lie (super)algebras
The notion of $T^{*}$-extension of Lie algebras was initially introduced by Bordemann [Bor]. A short explanation is as follows. Given a Lie algebra $\mathfrak{h}$ equipped with an $\mathfrak{h}^{*}$-valued bilinear map $\alpha$, a $T^{*}$-extension of $\mathfrak{h}$ is the Lie algebra structure on the vector space $\mathfrak{g}:=\mathfrak{h}\oplus\mathfrak{h}^{*}$ constructed by means of the form $\alpha$ that satisfies a certain cohomological requirement. The Lie algebra $\mathfrak{g}$ naturally admits a non-degenerate invariant symmetric bilinear form. Conversely, Bordemann showed in [Bor] that every finite-dimensional nilpotent Lie algebra $\mathfrak{h}$ with an invariant symmetric non-degenerate bilinear form “is” a suitable $T^{*}$-extension.
In particular, if $\mathfrak{h}$ is a real finite-dimensional Lie algebra belonging to a Lie group $G$, then the $T^{*}$-extension of $\mathfrak{h}$ is the Lie algebra of the cotangent bundle $T^{*}G$ of $G$, considered as a Lie group. The terminology and notation are justified by this differential geometric fact.
A superization of this construction was given in [BBB]. In particular, it was shown that every solvable Lie superalgebra with an even non-degenerate symmetric bilinear form is either isomorphic to a $T^{*}$-extension of a certain Lie superalgebra or to an ideal of codimension one of a certain $T^{*}$-extension.
@@
Lie algebras possessing both a non-degenerate invariant symmetric bilinear form and a non-degenerate 2-cocycle of its scalar cohomology, were studied in [BBM] and were called quadratic symplectic Lie algebras. If the ground field is algebraically closed, the authors showed that every quadratic symplectic Lie algebra is a $T^{*}$-extension of a Lie algebra which admits an invertible derivation. Necessary and sufficient conditions on $\mathfrak{h}$ and on the cocycle used in the construction of the $T^{*}$-extension are investigated to ensure that the extended algebra admits a skew-symmetric derivation and, hence, a symplectic structure. As a result, complex symplectic quadratic Lie algebras with dimensions less than or equal to 8 are classified. Some of these Lie algebras correspond to some Lie groups that admit a bi-invariant pseudo-Riemannian metric as well as a left-invariant symplectic form. These Lie groups are nilpotent, and their geometry is very rich, see [BBM, MR] for more details.
The notion of Lagrangian extension was superized in [BM] (see also [BBE] where the ground field is of characteristic 2). The construction is valid for arbitrary Lie superalgebras defined over a field of arbitrary characteristic provided the Lie superalgebra admit a torsion-free flat connection. Moreover, they have observed that there are two ways to perform this extension: they consider either $\mathfrak{g}:=\mathfrak{h}\oplus\mathfrak{h}^{*}$ or $\mathfrak{g}:=\mathfrak{h}\oplus\Pi(\mathfrak{h}^{*})$, where $\Pi$ is the change of parity functor. They call them $T^{*}$-extension and $\Pi T^{*}$-extensions of $\mathfrak{h}$, respectively. In the first case, the non-degenerate form is even, and in the second case, it is odd.
@@
One of the main objectives of this paper is to complete the list of examples among the 4-dimensional Lie superalgebras and explain how they can be obtained as Lagrangian extensions, see Section 4. According to the Backhouse classification, this investigation will classify all four-dimensional real Lie superalgebras that admit (or do not admit) Lagrangian extensions.
@@ 1.2. Left-symmetric structures
Left-symmetric algebras (also known in the literature as Vinberg-Koszul algebras) play an important role in various fields of mathematics and mathematical physics; for instance, rooted tree algebras, operad theory, vertex algebras, convex homogeneous cones, affine manifolds, left-invariant affine structures on Lie groups, to name just a few. For a thorough review, see [B3].
Over the real and complex numbers left-symmetric algebras are closely related to the theory of left-invariant affine structures over Lie groups; see, for example, [M, GS] (and references therein).
The investigation and classification of left-symmetric structure on Lie (super)algebras is an important subject on its own, as initiated in [S]. The problem, however, seems to be difficult to handle.
For reductive Lie algebras, Baues [Ba] proved that $\mathfrak{gl}(n)$ is the only reductive Lie algebra with one-dimensional center and a simple semisimple ideal which admits left- symmetric algebras over an algebraically closed ground field (see also [B1]). Moreover, all left-symmetric algebras on $\mathfrak{gl}(n)$ were classified.
Simple Lie algebras over complex numbers do not admit left-symmetric structures, see [H]. The existence of such structures is not always guaranteed, even for nilpotent Lie algebras. There are filiform nilpotent Lie groups of dimension $10\leq n\leq 13$ that do not admit any left-invariant affine structure, see [B, B3, BG]. Any filiform nilpotent Lie group of dimension $n\leq 9$ admits a left-invariant affine structure.
Nevertheless, left-symmetric algebra structures on both finite-dimensional and infinite-dimensional Lie algebras have been studied extensively, see [Bai, B2, YZ] as well as [CL, KCB, TB], respectively. Left-symmetric structures on the Virasoro superalgebra were studied in [KB].
In the super setting, it was proved in [DZ] that $\mathfrak{sl}(m+1|m)$ admits a left-symmetric superalgebra, and a full classification is offered in the case of $\mathfrak{sl}(2|1)$. Moreover, it was shown that $\mathfrak{sl}(m|1)$ does not admit a left-symmetric structure for $m\geq 3$.
Over a field of characteristic 2, all left-symmetric structures on a 2-dimensional superspace have been classified in [BBE].
@@
A second aim of this paper is to provide explicitly a left-symmetric structure on the 4-dimensional real Lie superalgebras. Surprisingly enough, all these Lie superalgebras admit such a structure, given explicitly in Section 5. Furthermore, we show that these structures are Novikov, except for the two Lie superalgebras $(D_{0}^{10})^{1}$ and $(D_{0}^{10})^{2}$. In addition, all these Lie superalgebras admit a Balinsky-Novikov structure.
Classifying all left-symmetric structures on each of these Lie superalgebras remains an open problem.
@@ 2. Main concepts and definitions
Throughout the text, $\mathbb{K}$ is a field of characteristic $p\not=2$.
Let $V=V_{\bar{0}}\oplus V_{\bar{1}}$ be a superspace defined over $\mathbb{K}$}. The parity of a homogeneous element $v\in V_{\bar{i}}$ is denoted by $|v|:=\bar{i}$. The element $v$ is called even if $v\in V_{\bar{0}}$ and odd if $v\in V_{\bar{1}}$. The superdimension of $V$ is $\mathrm{sdim}(V)=a+b\epsilon$, where $\epsilon^{2}=1$, and $a=\mathrm{dim}(V_{\bar{0}})$, $b=\mathrm{dim}(V_{\bar{1}})$. Usually, $\mathrm{sdim}(V)$ is shorthanded as $a\mid b$.
We will denote by $\Pi$ the functor that assigns to $V$ the superspace $\Pi(V)$ defined as the other copy of $V$ with the opposite parity of its homogeneous components. We have
@@
Let $V$ and $W$ be two superspaces defined over an arbitrary field $\mathbb{K}$}. Let $\mathcal{B}\in\text{Bil}(V,W)$ be a homogeneous bilinear form. The Gram matrix $B=(B_{ij})$ associated to $\mathcal{B}$} is given by the formula
$B_{ij}=(-1)^{|{\mathcal{B}}||v_{i}|}{\mathcal{B}}(v_{i},w_{j})\text{\penalty 10000\ \penalty 10000\ for the basis vectors $v_{i}\in V$ and $w_{j}\in W$.}$
[…]
From now on, let us assume that $V=W$.
In terms of the Gram matrix $B$ of $\mathcal{B}$}: the form $\mathcal{B}$} is symmetric if and only if
$u(B)=B,\;\text{ where $u(B)=\left(\begin{array}[]{cc}R^{t}&(-1)^{|\omega|}T^{t}\\[5.69054pt] (-1)^{|\omega|}S^{t}&-U^{t}\end{array}\right),$ for $B=\left(\begin{array}[]{cc}R&S\\[5.69054pt] T&U\end{array}\right)$.}$
Similarly, anti-symmetry of $\mathcal{B}$} means that u(B)=Bu(B)=-B.
Throughout the text, we denote by $\omega$ an anti-symmetric bilinear form.
An even (resp. odd) non-degenerate anti-symmetric bilinear form is called orthosymplectic (resp. periplectic).
Proposition 2.1 (Superdimension constraints).
If $V=V_{\bar{0}}\oplus V_{\bar{1}}$ is a finite-dimensional superspace such that $V_{\bar{1}}\not={0}$, equipped with a non-degenerate anti-symmetric bilinear form $\omega$, then:
  1. (i)
    If $|\omega|=\bar{0}$, then $\mathrm{dim}(V_{\bar{0}})$ is even.
  2. (ii)
    If $|\omega|=\bar{1}$, then $\mathrm{dim}(V_{\bar{0}})=\mathrm{dim}(V_{\bar{1}})$.
Proof.
See, [BM]
A Lie superalgebra is a superspace $\mathfrak{g}=\mathfrak{g}_{\bar{0}}\oplus\mathfrak{g}_{\bar{1}}$ over a field of characteristic $p\not=2$ together with a bilinear binary operator $\cdot,\cdot]:\mathfrak{g}\times\mathfrak{g}\rightarrow\mathfrak{g}$, called the Lie bracket that has the following properties: (for all homogeneous $x,y,z\in\mathfrak{g}$)
  1. (i)
    Super anti-commutativity: $x,y]=-(-1)^{|x||y|}[y,x]$;
  2. (ii)
    Super Jacobi identity: $x,[y,z]]=[[x,y],z]+(-1)^{|x||y|}[y,[x,z]]$ where $|x|={\bar{0}}$ if $x\in\mathfrak{g}_{\bar{0}}$ and $|x|={\bar{1}}$ if $x\in\mathfrak{g}_{\bar{1}}$;
  3. (iii)
    If $p=3$, one must add $[x,[x,x]]=0$ if $x\in\mathfrak{g}_{\bar{1}}$.
Let $\omega\in\mathrm{Bil}(\mathfrak{g},\mathfrak{g})$ be an anti-symmetric bilinear form on a Lie superalgebra $\mathfrak{g}$. The form $\omega$ is called closed if it satisfies
@@
for all homogeneous $x,y,z\in\mathfrak{g}_{\bar{i}}$. This means that $\omega$ is 2-cocycle with scalar values.
A Lie superalgebra $\mathfrak{g}$ is called quasi-Frobenius if it is equipped with a closed non-degenerate anti-symmetric bilinear form $\omega$. If $\omega$ is even (resp. odd) we call it orthosymplectic (resp. periplectic). We denote such a Lie superalgebra by $(\mathfrak{g},\omega)$. Recently, these Lie superalgebras were studied in [BBE, BE, BEM, BM] in the context of Lagrangian and double extensions.
@@ 2.1. Connections on Lie (super)algebras
Let $\nabla:\mathfrak{g}\times\mathfrak{g}\rightarrow\mathfrak{g}$ be a connection on $\mathfrak{g}$. For any homogeneous $x,y,z\in\mathfrak{g}_{\bar{i}}$, the torsion $T^{\nabla}$ of $\nabla$ is defined as
@@
2.2. Left-symmetric (super)algebras
@@
These superalgebras are Lie-admissible. The supercommutator $x,y]:=x\cdot y-(-1)^{|x||y|}y\cdot x$ defines then a Lie superalgebra structure on $L$ and we denote it by ${\mathfrak{g}}(L)$. The resulting Lie superalgebra is called the associated Lie superalgebra of the LSSA. Note that if $L$ admits an LSSA then the even part $L_{\bar{0}}$ also admits one. Associative superalgebras are all LSSAs.
In [H], Helmstetter showed that it is necessary to have $L,L]\subsetneq L$ if $L$ is a left-symmetric algebra defined over the field of complex numbers. Over a field of characteristic $p$, this is not true anymore. Burde showed in [B2] that classical Lie algebras over a field $p>3$ admit an LSA only in case $p$ divides $\mathrm{dim}(L)$. In the super setting and over a field of prime characteristic, there has not been much research done so far.
@@
Additionally, every quasi-Frobenius Lie superalgebra $({\mathfrak{g}},\omega)$ admits a left-symmetric structure given by
@@
The converse is not always true. Section 5 provides several examples illustrating this.
Certain subclasses of LSSAs are also of particular interest. An LSSA with supercommutative right multiplication is called a Novikov superalgebra (cf. [BN, Xu]). Explicitly, a Novikov superalgebra $(L,\cdot)$ is an LSSA with an additional condition
@@
    Left-symmetry and commutativity:
@@
[…]
@@
[…]